Find the Jacobian of the transformation.

x = {\color{red}} e^{{\color{red}2} s + {\color{red}5} t} \mbox{, } y = e^{{\color{red}2} s - {\color{red}5} t} $
Answer #1

| 2S+st 으e . 25-6 = 2 au Nv) 1 년 0 0) 0/ aV OK ? ost) 1 | 꿇 쭖 2S+SE. tsts - 음 (255t) - 2.tes + st) 을 25 = 512sty 3년 - 2 (23-s25+57 +25-$ t a 10.e $+$+25-5€. = -0.e = -10.645-10eus -20.eus